# On Finding Small Hyper-Gradients in Bilevel Optimization: Hardness Results and Improved Analysis

### v4

@@ Appendix B Future Works
(e). Pick $y_{1},y_{2}$ as Equation 13. Similarly, there exist $y_{1}^{\prime}\in Y^{*}(x_{1})$ and $y_{2}^{\prime}\in Y^{*}(x_{2})$ such that
@@
Appendix F Proof of Theorem 3.2
Our hard instance is based on the following convex zero-chain [42], Section 2.1.2.
Definition F.1 (Worse-Case Zero-Chain).
@@
In bilevel problems, it is crucial to find a point $y$ that is close to $Y^{*}(x)$, instead of just achieving a small optimality gap $g(x,y)-g^{*}(x)$. However, it is difficult for any first-order algorithms to “locate” the minimizers of the function class in Definition F.1. Below, we formalize this observation into a rigorous statement.
See 3.2
Proof.
@@
where $h_{q}(y)$ follows Definition F.1 and $r(y)=\sum_{j=q/2+1}^{q}\psi(y_{[j]})$, where $\psi:{\mathbb{R}}\rightarrow{\mathbb{R}}$ is
@@
Appendix G Proof of Theorem 4
First of all, we recall some useful lemmas for PL conditions [26].
Lemma G.1.
@@ Appendix G Proof of Theorem 4
The above two lemmas appear in Theorem 2 [26]. Based on these results, we can prove the following lemma.
See 4.1
Proof.
By Lemma G.1, for any $y_{1}\in Y_{\sigma_{1}}^{*}(x_{1})$, there exists some $y_{2}\in Y_{\sigma_{2}}^{*}(x_{2})$ such that
@@
By symmetry, for any $y_{2}\in Y_{\sigma_{2}}^{*}(x_{2})$, there also exists $y_{1}\in Y_{\sigma_{1}}^{*}(x_{1})$ such that that the above inequality holds for $y_{2},y_{2}$. This proves the Pompeiu–Hausdorff continuity. ∎
See 4.2
Proof.
It follows the generalized Danskin’s theorem proved in [49]. See also Lemma A.2 in [29]. ∎
See 4.3
Proof.
The proof follows Theorem 3.8 [29]. The only difference is that Theorem 3.8 [29] states $\|\nabla\varphi_{\sigma}(x)-\varphi(x)\|={\mathcal{O}}(\sigma\ell\kappa^{5})$. But the additional $\kappa^{2}$ dependency comes from the perturbation in the multiplier ${\rm d}\lambda/{\rm d}\sigma\asymp\kappa^{2}$. Since we only consider the unconstrained case, there is no need to consider the effect of ${\rm d}\lambda/{\rm d}\sigma$, and we can improve the bound to ${\mathcal{O}}(\sigma\ell\kappa^{3})$. ∎
We also recall some technical lemmas from [29].
Lemma G.3.
@@
Proof.
See Proposition 3.1 [29]. ∎
We remark that Proposition 6 [2] also presents a similar argument as the above lemma for Morse-Bott functions.
Lemma G.4.
@@
Finally, we plug in the explicit form of $\dfrac{\partial^{2}}{\partial x\partial\sigma}l(x,\sigma)$ by Theorem 3.2 [29].
@@
Similarly, if there exists some matrix $Q$ such that $U=QA$, then $U=UAA^{\dagger}$. Combining these two identities completes the proof. ∎
@@
On the one hand, Lemma G.2 indicates
@@
Using $\alpha_{r}=o(t)$ and letting $t\rightarrow 0$, we conclude that $\lambda_{r}\geq\mu$. ∎
Lemma 4.3 only claims the existence of $\nabla\varphi(x)$. Below, we give the explicit form of $\nabla\varphi(x)$.
Lemma G.7.
@@
Proof.
Let $H(x)$ be the right-hand side of Equation 15. Below, we show that $\nabla\varphi(x)=H(x)$.
Recall Lemma G.4 that for any $\sigma^{\prime}\geq 0$ there exists some $\sigma\in[0,\sigma^{\prime}]$ such that
@@
for any $y_{\sigma}^{*}(x)\in Y_{\sigma}^{*}(x)$. Then for any $y^{*}(x)\in Y^{*}(x)$, by Lemma 4.1, there exists $y_{\sigma}^{*}(x)\in Y_{\sigma}^{*}(x)$ such that $\|y_{\sigma}^{*}(x)-y^{*}(x)\|\leq C_{f}\sigma/\mu$. Then by Lemma G.3 and Lemma G.5, we have
@@
We then use Lemma G.6 to have a further upper bound as
@@
See 4.4
Proof.
Invoking Lemma 4.1, there exists $y_{1}\in Y^{*}(x_{1})$ and $y_{2}\in Y^{*}(x_{2})$ such that $\|y_{1}-y_{2}\|\leq L_{g}/\mu$. Then by Lemma G.3 and Lemma G.5, we have
@@
Further invoking Lemma G.6,
@@
Then by Lemma G.8, we have
@@
By Young’s inequality and Lemma 4.1,
@@
Plug into Equation 19, which, in conjunction with Equation 17, yields that
@@
By Lemma 4.1 we know that for any
@@
Appendix H Proof of Theorem 4.2
Lemma H.1.
@@
We conclude the proof by using Equation 16.
See 4.2
Proof.
Define $\delta_{t}:={\mathbb{E}}\left[{\rm dist}^{2}\left(y_{t}^{0},Y_{\sigma}^{*}(x_{t})\right)+{\rm dist}^{2}\left(z_{t}^{0},Y^{*}(x_{t})\right)\right]$. By Equation 18, letting
@@
ensures ${\mathbb{E}}\|\hat{\nabla}\varphi(x_{t})-\nabla\varphi_{\sigma}(x_{t})\|^{2}\leq{\mathcal{O}}(\epsilon)$. Then telescoping over Equation 17 shows that one can find an $\epsilon$-stationary point of $\varphi(x)$ with $T={\mathcal{O}}(\epsilon^{-2})$ outer-loop iterations. By Lemma H.1, it suffices to set the parameters in the inner loop as Equation 8. But $K_{t}$ requires the knowledge of $\delta_{t}$. Next, we bound $\delta_{t}$ via a recursion which allows us to get rid of the prior knowledge of $\delta_{t}$. By Lemma H.1 and Lemma 4.1,
@@
we can get Equation 9. Telescoping over $t$,
@@
where on the right-hand side we know $\sum_{t=0}^{T-1}\|x_{t+1}-x_{t}\|$ must also be bounded by Equation 17.
@@
And the total number of stochastic oracle calls is $B$ times the above bound for iterations. ∎

### v6

@@ Appendix D Proof of Theorem 1
(e). Pick $y_{1},y_{2}$ as Equation (13). Similarly, there exist $y_{1}^{\prime}\in Y^{*}(x_{1})$ and $y_{2}^{\prime}\in Y^{*}(x_{2})$ such that
@@
Appendix E Proof of Theorem 2
Our hard instance is based on the following convex zero-chain. The following function is very similar to the worst-case function given by (Nesterov, 2018), Section 2.1.2. The only difference is that the function by Nesterov (2018) has an additional term $z_{[q]}^{2}/8$.
Definition E.1 (Worse-Case Zero-Chain)
@@ Appendix E Proof of Theorem 2
Proof E.1.
We prove each property one by one. It can easily be seen that $h_{q}(z)\geq 0$ for all $z\in{\mathbb{R}}^{q}$ and the equality holds if and only if $z={\bf{1}}$. This proves property a. Further, note that $h_{q}(z)$ is quadratic with Hessian given by
$A=\frac{1}{4}\begin{bmatrix}2&-1&&&\\ -1&2&-1&\\ &-1&2&-1&\\ &&\ddots&\ddots&\ddots\\ &&&-1&1\end{bmatrix}.$
As $A$ is diagonally dominant, we know that $A\succeq O$. This proves property b. For any $v\in{\mathbb{R}}^{q}$,
$v^{\top}Av$ $=\frac{1}{4}\left[v_{[1]}^{2}+\sum_{j=1}^{q-1}(v_{[j]}-v_{[j+1]})^{2}\right]$
$\leq\frac{1}{4}\left[v_{[1]}^{2}+\sum_{j=1}^{q-1}(v_{[j]}-v_{[j+1]})^{2}+v_{[q]}^{2}\right]$
$\leq\frac{1}{4}\left[v_{[1]}^{2}+\sum_{j=1}^{q-1}2(v_{[j]}^{2}+v_{[j+1]}^{2})+v_{[q]}^{2}\right]$
$\leq\sum_{j=1}^{q}v_{[j]}^{2}=\|v\|^{2}.$
This proves property c. Finally property d holds since $A$ is tridiagonal.
In bilevel problems, it is crucial to find a point $y$ that is close to $Y^{*}(x)$, instead of just achieving a small optimality gap $g(x,y)-g^{*}(x)$. However, it is difficult for any first-order algorithms to “locate” the minimizers of the function class in Definition E.1. Below, we formalize this observation into a rigorous statement.
See 2
Proof E.2.
@@
where $h_{q}(y)$ follows Definition E.1 and $r(y)=\sum_{j=q/2+1}^{q}\psi(y_{[j]})$, where $\psi:{\mathbb{R}}\rightarrow{\mathbb{R}}$ is
@@
Appendix F Proof of Theorem 4
First of all, we recall some useful lemmas for PL conditions (Karimi et al., 2016).
Lemma 1.
@@ Appendix F Proof of Theorem 4
The above two lemmas appear in Theorem 2 (Karimi et al., 2016). Based on these results, we can prove the following lemma.
See 2
Proof F.1.
By Lemma 1, for any $y_{1}\in Y_{\sigma_{1}}^{*}(x_{1})$, there exists some $y_{2}\in Y_{\sigma_{2}}^{*}(x_{2})$ such that
@@
By symmetry, for any $y_{2}\in Y_{\sigma_{2}}^{*}(x_{2})$, there also exists $y_{1}\in Y_{\sigma_{1}}^{*}(x_{1})$ such that that the above inequality holds for $y_{2},y_{2}$. This proves the Pompeiu–Hausdorff continuity.
See 3
Proof F.2.
It follows the generalized Danskin’s theorem proved in Shen and Chen (2023). See also Lemma A.2 in Kwon et al. (2024b).
See 4
Proof F.3.
The proof follows Theorem 3.8 Kwon et al. (2024b). The only difference is that Theorem 3.8 Kwon et al. (2024b) states $\|\nabla\varphi_{\sigma}(x)-\varphi(x)\|={\mathcal{O}}(\sigma\ell\kappa^{5})$. But the additional $\kappa^{2}$ dependency comes from the perturbation in the multiplier ${\rm d}\lambda/{\rm d}\sigma\asymp\kappa^{2}$. Since we only consider the unconstrained case, there is no need to consider the effect of ${\rm d}\lambda/{\rm d}\sigma$, and we can improve the bound to ${\mathcal{O}}(\sigma\ell\kappa^{3})$.
We also recall some technical lemmas from Kwon et al. (2024b).
Lemma 3.
@@
Proof F.4.
See Proposition 3.1 (Kwon et al., 2024b).
We remark that Proposition 6 (Arbel and Mairal, 2022) also presents a similar argument as the above lemma for Morse-Bott functions.
Lemma 4.
@@
Finally, we plug in the explicit form of $\dfrac{\partial^{2}}{\partial x\partial\sigma}l(x,\sigma)$ by Theorem 3.2 (Kwon et al., 2024b).
@@
Similarly, if there exists some matrix $Q$ such that $U=QA$, then $U=UAA^{\dagger}$. Combining these two identities completes the proof.
@@
On the one hand, Lemma 2 indicates
@@
Using $\alpha_{r}=o(t)$ and letting $t\rightarrow 0$, we conclude that $\lambda_{r}\geq\mu$.
Lemma 4 only claims the existence of $\nabla\varphi(x)$. Below, we give the explicit form of $\nabla\varphi(x)$.
Lemma 7.
@@
Proof F.8.
Let $H(x)$ be the right-hand side of Equation (15). Below, we show that $\nabla\varphi(x)=H(x)$.
Recall Lemma 4 that for any $\sigma^{\prime}\geq 0$ there exists some $\sigma\in[0,\sigma^{\prime}]$ such that
@@
for any $y_{\sigma}^{*}(x)\in Y_{\sigma}^{*}(x)$. Then for any $y^{*}(x)\in Y^{*}(x)$, by Lemma 2, there exists $y_{\sigma}^{*}(x)\in Y_{\sigma}^{*}(x)$ such that $\|y_{\sigma}^{*}(x)-y^{*}(x)\|\leq C_{f}\sigma/\mu$. Then by Lemma 3 and Lemma 5, we have
@@
We then use Lemma 6 to have a further upper bound as
@@
See 5
Proof F.9.
Invoking Lemma 2, there exists $y_{1}\in Y^{*}(x_{1})$ and $y_{2}\in Y^{*}(x_{2})$ such that $\|y_{1}-y_{2}\|\leq L_{g}/\mu$. Then by Lemma 3 and Lemma 5, we have
@@
Further invoking Lemma 6,
@@
Then by Lemma 8, we have
@@
By Young’s inequality and Lemma 2,
@@
Plug into Equation (19), which, in conjunction with Equation (17), yields that
@@
By Lemma 2 we know that for any
@@
Appendix G Proof of Theorem 7
Lemma 1.
@@ Appendix G Proof of Theorem 7
We conclude the proof by using Equation (16).
See 7
Proof G.2.
Define $\delta_{t}:={\mathbb{E}}\left[{\rm dist}^{2}\left(y_{t}^{0},Y_{\sigma}^{*}(x_{t})\right)+{\rm dist}^{2}\left(z_{t}^{0},Y^{*}(x_{t})\right)\right]$. By Equation (18), letting
@@
ensures ${\mathbb{E}}\|\hat{\nabla}\varphi(x_{t})-\nabla\varphi_{\sigma}(x_{t})\|^{2}\leq{\mathcal{O}}(\epsilon)$. Then telescoping over Equation (17) shows that one can find an $\epsilon$-stationary point of $\varphi(x)$ with $T={\mathcal{O}}(\epsilon^{-2})$ outer-loop iterations. By Lemma 1, it suffices to set the parameters in the inner loop as Equation (8). But $K_{t}$ requires the knowledge of $\delta_{t}$. Next, we bound $\delta_{t}$ via a recursion which allows us to get rid of the prior knowledge of $\delta_{t}$. By Lemma 1 and Lemma 2,
@@
we can get Equation (9). Telescoping over $t$,
@@
where on the right-hand side we know $\sum_{t=0}^{T-1}\|x_{t+1}-x_{t}\|$ must also be bounded by Equation (17).
@@
And the total number of stochastic oracle calls is $B$ times the above bound for iterations.